OpenStudy (goldphenoix):

The median of (1/6), (2/5), (3/4), (4/3), (5/2), and (6/1) is A. 1 B. (669/360) C. (7/12) D. (25/24)

OpenStudy (anonymous):

are those supposed to be points?

OpenStudy (goldphenoix):

What do you mean?

OpenStudy (whpalmer4):

Line them up in order from smallest to largest. You've got an even number of data points, so the median is the mean of the middle two values. \[\frac{1}{2}(\frac{3}{4} + \frac{4}{3}) =\]

OpenStudy (goldphenoix):

Wow... Why am I getting mixed up with multiplication and addition. T~T. I see, so the answer is D. (25/24)

OpenStudy (whpalmer4):

If you had an odd number, it would just be the middle value, no calculation required.

OpenStudy (whpalmer4):

D is correct.

OpenStudy (goldphenoix):

I thought you had to multiply 3/4 and 4/3 then divide it by 2. T~T

OpenStudy (anonymous):

oh.. statistic median. i was thinking of graphical median. wew

OpenStudy (anonymous):

those parentheses are deceiving

OpenStudy (goldphenoix):

Yeah, sorry.

OpenStudy (goldphenoix):

those were supposed to be fractions.

OpenStudy (whpalmer4):

better more parentheses than needed than the usual openstudy case which is not nearly enough!